The length of a rectangle is 3 times its width. If the perimeter of the rectangle is 32 cm, find its length and width.
Answer by shweta(56)

Let 'l' be the length of the rectangle and 'w' be the width of the rectangle
Given:l=3w and Perimeter,P=32 cm
P= 2*(l+w)
Substitute the given value for l
32 = 2(3w+w)
32/2 = 4w
4w = 16
w= 16/4
w= 4
l= 3*w
l= 3*4
l= 12
Length of a rectangle is 12cm and width is 4cm